OpenStudy (anonymous):

FIND AN EQUATION OF THE TANGENT LINE TO THE GRAPH OF f AT THE GIVEN POINT. (X+1)/(X-1), {1/2,-3}

OpenStudy (anonymous):

take the derivative, plug in 1/2

OpenStudy (anonymous):

AND F(X)=(1+SINX)/(1-SINX), {(PIE,1}

OpenStudy (anonymous):

derivative of \[\frac{x+1}{x-1}\] is \[-\frac{2}{(x-1)^2}\] replace x by 1/2 to get your slope

OpenStudy (anonymous):

second problem same as first. derivative is \[\frac{2\cos(x)}{(1-\sin(x))^2}\]

OpenStudy (anonymous):

FOR FIRST DERIVATIVE WHY DID I GET -X/{2(X-1)^2???

OpenStudy (anonymous):

ONCE I FIND THE SLOPE THEN WHAT??

OpenStudy (anonymous):

Satellite's derivative is correct. To avoid mistakes in future, you should not simplify the derivative: there is no need to break it down since you are just inputting values for x; that only increases chances to make a mistake, and in exam waste time.

OpenStudy (anonymous):

Once you find slope, equation is y=[f'(1/2)][x-(1/2)] - 3